HelpMePlz:

suppose y varies directly with x. If x = 18, when y = 10 what is the value of x when y = 12

Champion:

y varies directly with x so, y = kx given that x= 18, y= 10, can you plug these in y = kx and get the value of k ??

HelpMePlz:

i got k = 11.6? but doesnt work.

Champion:

because we need the value of 'x' ! not k

HelpMePlz:

still confused

Champion:

y = kx 10 = 18k k = 10/18. y = kx 12 = (10/18) x x = 12*18/10 = ..?

HelpMePlz:

ohhhh!! thank you!!!

Champion:

welcome ^_^